Grade 11Trigonometry

Solve the eqn:
3tan(2x)-4tan(3x)=tan2(3x)tan(2x)
Get the general solution of this eqn

Ishaan
7 Years agoGrade 11

1 Answer

Deepak Kumar Shringi
7 Years ago
3 tan 2x - 4 tan 3x = tan² 3x * tan 2x

3 (tan 2x - tan 3x) = tan 3x + tan² 3x * tan 2x
= tan 3x (1 + tan 3x * tan 2x)

3 (tan 2x - tan 3x)/(1 + tan 3x * tan 2x) = tan 3x

3 * (- tan x) = tan 3x
-3tanx = (3 tan x - tan³ x) / (1 - 3 tan² x)

3 - tan² x = - 3 + 9 tan² x

tan x = + √3 /√5 or - √3/√5